OpenStudy (anonymous):

Compensation, solve the equation below 221+335=? Remember I need to compensate!

OpenStudy (anonymous):

556? is that it;?

OpenStudy (anonymous):

well I need to compensate like subtract 3 for example and add it to the other number then subtract or add it mentally

OpenStudy (anonymous):

...

OpenStudy (anonymous):

Well, do get the compensation?

OpenStudy (anonymous):

oww ..

OpenStudy (anonymous):

because i just don't understand how to compensate

OpenStudy (anonymous):

just a sec

OpenStudy (anonymous):

you have to make both number to be in 10th like ex: 98+100 = 200 (98 +2) + (102-2) =100+100=200 so the numbers from both sides have to be the same like +2 or 02 it doesnt metter cause the answer is going to be the same .. but when you add that number and make it like a whole with a 0 in the end it's easier to add... do you understand?

OpenStudy (anonymous):

-2*

OpenStudy (anonymous):

Lifesaver thanks I really needed that!

OpenStudy (anonymous):

you welcome :_)